Circular No. 7466

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1859+226 J. E. McClintock, Harvard-Smithsonian Center for Astrophysics (CfA);

R. A. Remillard, Massachusetts Institute of Technology; and W. A. Heindl and J. A. Tomsick, University of California at San Diego, report: “We obtained 25 hr of photometry for this x-ray nova on July 3–7 UT, using the 1.2-m telescope at the Fred L. Whipple Observatory. During 6 hr of continuous monitoring on July 3, the source was constant in intensity and relatively faint (V = 19.2). On the following four nights, the source was brighter by ∼ 0.6 mag and varied nightly by up to 0.2 mag. We searched these four nights of data for a periodic modulation, using the technique of Stellingwerf (1978, Ap.J. 224, 953). The deepest minimum in the variance statistic occurs at 18.72 ± 0.48 hr, which we identify as the orbital period. We find no evidence for the reported 6.73-hr period (IAUC 7303). However, the reported 9.15-hr period (IAUC 7388) may be a subharmonic of the 18.7hr period reported herein. The light curve folded on the 18.7-hr period resembles that of GRO J1655–40 as it approached quiescence (Bailyn et al. 1995, Nature 378, 157). There is a primary minimum with a depth of ∼ 0.2 mag that is probably due to the partial eclipse of the accretion disk by the secondary star. There is also a shallow secondary minimum that is displaced by ∼ 0.5 in phase from the primary minimum; this feature is probably due to the eclipse of the secondary star by the disk. The time of the primary minimum is HJD 2451732.81 ± 0.03. There is a second deep mimimum in the periodogram at twice 18.7 hr; however, the complexity of features in the folded light curve argues that this is not the orbital period.”

M. Garcia and J. McClintock, CfA; P. Callanan, University College, Cork; and A. V. Filippenko, University of California at Berkeley, report, following IAUC 7463, that they observed the optical counterpart to this soft x-ray transient (cf. IAUC 7451, 7456) with the Cerro Tololo 0.9-m telescope on July 23.1, 24.1, and 25.1 UT, yielding the following R magnitudes, respectively: 20.8 ± 0.3, 20.2 ± 0.1, and 19.8 ± 0.1. Seeing was 2′′.7, 1′′.5, and 1′′.4 on the respective nights, with thin clouds on the first and clear skies on the next two. Thus this object appears to be undergoing yet another optical brightening. Observations on the Keck-II telescope also show a brightening in the K band between July 22 and 24.

SUPERNOVA 2000cr IN NGC 5395 CCD R magnitudes by K. Hornoch, Lelekovice, Czech Republic (0.35-

m reflector): July 6.943 UT, 16.5; 12.910, 16.4.

2000 July 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green